# Tradeoffs Between Alignment and Helpfulness in Language Models with Representation Engineering

### v4

@@ Abstract
Language model alignment has become an important component of AI safety, allowing safe interactions between humans and language models, by enhancing desired behaviors and inhibiting undesired ones. It is often done by tuning the model or inserting preset aligning prompts. Recently, representation engineering, a method which alters the model’s behavior via changing its representations post-training, was shown to be effective in aligning LLMs (Zou et al., 2023a). Representation engineering yields gains in alignment oriented tasks such as resistance to adversarial attacks and reduction of social biases, but was also shown to cause a decrease in the ability of the model to perform basic tasks. In this paper we study the tradeoff between the increase in alignment and decrease in helpfulness of the model. We propose a theoretical framework which provides bounds for these two quantities, and demonstrate their relevance empirically. First, we find that under the conditions of our framework, alignment can be guaranteed with representation engineering, and at the same time that helpfulness is harmed in the process. Second, we show that helpfulness is harmed quadratically with the norm of the representation engineering vector, while the alignment increases linearly with it, indicating a regime in which it is efficient to use representation engineering. We validate our findings empirically, and chart the boundaries to the usefulness of representation engineering for alignment. Our code is available at https://github.com/dorin133/REPE_alignment_helpfulness_tradeoff.
@@ 1 Introduction
Advancements in large language model (LLM) development over the last few years have given LLMs a variety of abilities that allow them to serve as general purpose assistants in a wide range of tasks, such as broad-scoped question answering, writing assistance, teaching, and more  (Radford et al., 2019; Devlin et al., 2019; Brown et al., 2020; Schulman et al., 2023; OpenAI, 2023; Bubeck et al., 2023; Nori et al., 2023; West, 2023; Park et al., 2023a). The vast use of LLMs for such purposes has raised concerns due to the harm they can cause their users, such as serving fake information (Lin et al., 2022; Weidinger et al., 2022), behaving offensively, feeding social biases (Hutchinson et al., 2020; Venkit et al., 2022; Weidinger et al., 2022), or encouraging problematic behaviors by users Roose (2023); Atillah (2023). Alignment is often the term given for the process of removing these undesired behaviors (Yudkowsky, 2001; Taylor et al., 2016; Amodei et al., 2016; Shalev-Shwartz et al., 2020; Hendrycks et al., 2021; Pan et al., 2022; Ngo, 2022).
@@
There are several different approaches to performing alignment in LLMs, such as including aligning prompts Askell et al. (2021); Rae et al. (2021) which was shown to improve alignment and decrease toxicity in LLMs, and the procedure of reinforcement learning from human feedback (RLHF) which trains language models to be helpful and harmless (Bai et al., 2022). Though effective to an extent, these approaches are still dangerously frail, as several works have shown that adversarial prompts can trigger negative behaviors in LLMs  Wallace et al. (2019); Yu & Sagae (2021); Xu et al. (2021); Subhash (2023); Zou et al. (2023b). The work of Wolf et al. (2023) provides a theoretical framework which shows that frozen LLMs can be misaligned with sufficiently long prompts.
Recently, a new method for alignment has been proposed by Zou et al. (2023a), which controls the model at the internal representation level by adding tailored vectors to the hidden layers’ representations. This is done by extracting directions in the model’s latent space that connect contrasting behaviors, and then injecting vectors at inference time in order to steer away from undesired behaviors and towards desired ones. Typically, the vectors are all prepared with norm $1$, and they are multiplied by a coefficient to tune the strength of the steering, but there is a trade-off – when the parameter is too high performance tends to degrade. Zou et al. (2023a) demonstrated experimentally that the procedure can significantly improve alignment, e.g., in resistance to adversarial attacks, with reduction from 50% success of adversarial attacks to less than 15%, and truthfulness enhancement, with a relative increase of over 50%, though at the cost of somewhat reducing the helpfulness of the model. Similar methods have also been used by Jorgensen et al. (2023); Leong et al. (2023); Liu et al. (2023); Turner et al. (2023) to improve alignment and reduce toxicity.
@@
Since then, there has been an increasing body of work using this method. Wang et al. (2024b) use extracted safety vectors for inference time alignment for harmlessness, reducing jailbreaking success rate from over 30% with prompting and over 10% in supervised fine tuning to below one percent. Wang et al. (2024a) uses a method of editing model parameters that maximize the difference between toxic and untoxic responses to detoxify it. Wei et al. (2024) find sparse regions in parameter space that affect alignment brittleness, to be removed for better alignment. Marks et al. (2024) interpret causal graphs in language models and edit them to improve behaviors. van der Weij et al. (2024) extend activation steering to multiple behaviors. To improve low rank finetuning, Wu et al. (2024) utilize a procedure of tuning representations directly to substantially reduce the trainable parameters of finetuning compared to LoRA. Xu et al. (2024); Li et al. (2024) use concept activation vectors to jailbreak, they also observe that concepts that activate different behaviors are linearly separable. Zhang et al. (2024) remove hallucinations by editing truthfulness concepts. Additionally, the method scales to SOTA models, such as Sonnet’s Claude Templeton (2024), using a similar method of sparse auto encoders, which extracts interpretable features from the model that can be used to manipulate the model through steering.
There are also known limitations to editing representations. Limitations of model editing methods for social debiasing are studied in Yan et al. (2024), and the work of Elazar et al. (2021) empirically demonstrates how projecting out supervised linear probe directions via iterative nullspace projection can reduce performance on selected tasks. Theoretical works on the subject show that in context learning is equivalent to inserting a query with a task vector Hendel et al. (2023) and the vectors that connect representations of token pairs are parallel when the semantic difference between the pairs is similar (the linear representation hypothesis) Park et al. (2023b).
@@
Understanding the tradeoff between model helpfulness and alignment is important for designing safe yet useful LLM systems. Previous empirical works have shown tradeoffs between quality and diversity and between helpfulness and safety in LLMs due to instruct finetuning (Florian et al., 2024; Bianchi et al., 2023; Röttger et al., 2023), and reduction in performance due to watermarking (Ajith et al., 2023). In this work we aim to shed light on the benefits and limitations of representation engineering for LLM alignment, i.e., how much does alignment improve with this method and what is the cost in terms of the model’s abilities. We approach this question theoretically at first, and then provide empirical evidence for the validity of our theory.
@@
Figure 1: Effect of representation engineering on helpfulness and alignment. Our main results show that alignment can improve at the cost of helpfulness. Moreover, we show that for small representation engineering norms the helpfulness decreases quadratically while the alignment increase is linear, so there is a regime in which representation engineering can be cost-effective.
@@
In sections 2 and 3, we set up our theoretical framework and present our theoretical results respectively. We find that representation engineering increases alignment linearly with the representation engineering vector norm (theorem 1), while the helpfulness of the model, defined as the probability of answering a query correctly, decreases quadratically with the representation engineering vector norm (theorem 2). Consequently, alignment can be guaranteed with large enough representation injections (corollary 1), though at the cost of significantly reducing the model’s helpfulness, e.g.multiple choice question answering reduced to random guessing (corollary 2). Conversely, this means that when injecting vectors of small norms, the improvement of alignment is initially faster than the decrease in helpfulness, possibly indicating that there is a regime where representation engineering is more effective, allowing for inference time alignment while maintaining the model’s helpful capabilities. See figure 1 for an illustration of this intuition.
@@
In section 4 we explore the validity of our assumptions and results in an experimental setting: We calculate alignment, as defined by the theoretical framework, as a function of representation engineered vector norms corresponding to the desired behaviors and find that it increases as predicted by theorem 1. This is done by aligning with representation engineering an unaligned (pretrained) model with respect to desired behaviors (“harmless", “not-racist"), and misaligning an aligned (RLHF) model to undesired behaviors (“harmful", “racist"). Then, we calculate the helpfulness of the model, quantified by its question answering abilities over different knowledge domains and coding capabilities, with the same aligning vectors, and find that the decay with increased vector norm described in theorem 2 is manifested. Furthermore, injecting large norms of these vectors leads to significant reduction of helpfulness, such as performance on multiple-choice questions that is equivalent to random guessing (corollary 2). Together, the results correspond to the intuitive illustration in fig. 1.
@@ 2 Preliminaries
We denote $P_{\theta}(\cdot|s)$ as the model’s next token probability distribution with the parameters $\theta$ to the prompt $s$. The model is composed of $L$ layers, $r_{\theta}^{l}$ is the $l$’th hidden state representation of the model. The next token prediction of a model is parametrized as:
@@
Where $r^{(L)}_{\theta}(s)$ is the final hidden layer’s representation of the prompt $s$ and $U$ is a matrix from the hidden state to a vocabulary of tokens. This is an accurate parametrization for state-of-the-art LLMs.
Parameterizing representation engineering is done by modifying each layer’s hidden state via adding a corresponding engineered vector: Denote a representation engineered model $P_{\theta}$ of $L$ layers, with a set of engineered representations $R_{e}=(r_{e}^{(l=1)},...,r_{e}^{(l=L-1)})$ by $P_{\theta,r_{e}}$, which is applied by adding to each hidden state the corresponding engineered vector:
@@
Note that here $l<L$, as used in Zou et al. (2023a). Additionally, we follow existing methods for representation engineering and provide a uniform norm for all the injected vectors $|r_{e}^{(l)}|=|r_{e}|$. The vectors are initially prepared with norm $1$, and when injected to the model, they are multiplied by the coefficient $r_{e}$ which can be positive or negative, to tune the steering strength and direction. For layers that are not injected, $|r_{e}^{(l)}|=0$.
To quantify alignment, we use the behavior expectation definition of alignment as in Wolf et al. (2023). We will use a binary scoring function, with labels ±$\pm 1$ for aligned/misaligned answers. The results can be extended to more complex behavior scoring functions over $[-1,+1]$, to yield qualitatively similar results, as discussed appendix H:
@@
Notice that while $B$ is a binary function, the behavior expectation is in the range $[-1,+1]$, reflecting cases where an aligned response is required and unaligned responses must be filtered. In theorem 1 we will prove that representation engineering is an effective alignment method by lower bounding the behavior expectation. Notice that high probability of outputting a positive/negative response gives a positive/negative contribution to the behavior expectation, thus the sign and absolute value of behavior expectation is a good measure for the alignment of a model.
The model’s helpfulness can be quantified as its ability to produce useful answers to user’s queries (knowledge questions, code generation, summarization, etc.). In order to theoretically analyze helpfulness, we focus on queries where correctness can be defined, such as knowledge based question answering on various domains (see figure 1 for an example), or producing code to solve a problem. This can be measured as the likelihood of outputting a correct answer to a query:
@@
Where $P_{\theta}(a_{correct}|q)$ is the model’s probability of outputting the correct answer $a$ to the query $q$. By this definition, the helpfulness is in the range $[0,1]$, and motivation behind it is to quantify the general capabilities of the model when engineered representations are injected into it. For queries where correctness is not defined, the bounds we derive are expected to still be meaningful as they can also describe the rate of the model’s deviation from its original distribution due to representation engineering, as will be explained in the next section.
Ideally, a model that interacts with a user should be both aligned and helpful, meaning its response is appropriate w.r.t. a desired behavior (i.e., positive behavior expectation) and also useful (i.e., high probability of giving a correct answer to the query). In the next section, we will provide results on alignment and helpfulness under the use of representation engineering, based on the model’s next token prediction, which provides simple analytical forms for alignment and helpfulness. In appendix I, we extend the results for multi-token answers, which yields qualitatively similar results, with somewhat more complex form.
@@ 3 Main Results
We will show that representation engineering improves alignment, but harms helpfulness. Theorem 1 shows that behavior expectation is bounded from below by a hyperbolic tangent function, such that it approaches $+1$ for increasing size of injected vectors and increases linearly within a bounded range. This in principle allows to sample an aligned response for any adversarial attack (corollary 1), demonstrating the power of representation engineering as an alignment technique. Theorem 2 shows that the helpfulness is maximized in the vicinity of norm zero injected vectors (i.e., no representation engineering) and in corollary 2 that as the norm is increased, helpfulness decays to random guessing. The assumptions used to prove the theorems are presented formally in appendix A.
The following statement quantifies how alignment is improved by representation engineering. It assumes the injected representations in all layers accumulate to a change in the last hidden layer representation that classifies positive and negative behavior answers to the query, as depicted in figure 2a. This condition was chosen due to the popular choice in representation engineering to use injected representations $\{r_{e}^{(l)}\}$, that are themselves classifiers for positive and negative representations on the intermediate layers. This is because they are learned from contrasting positive and negative behavior representations for different queries, such as mean centering, $r_{e}^{(l)}=\mathbb{E}_{good,bad}[r_{good}^{(l)}-r_{bad}^{(l)}]$ (Jorgensen et al. (2023)), or PCA, $r_{e}^{(l)}=\text{argmax}_{v:||v||=1}[\mathbb{E}_{good,bad}|\langle v,r_{good}^{(l)}-r_{bad}^{(l)}\rangle|^{2}]$ (Zou et al. (2023a)), such that they form linear classifiers for the intermediate layers due to the positive/negative inner product with positive/negative answer representations. Notably, in Xu et al. (2024) it is shown empirically that such concept classes in latent space are linearly separable. We discuss this assumption further in A and provide empirical evidence. Furthermore, the classification condition can be softened to an imperfect classifier, as discussed in appendix A and shown in appendix in F, to yield similar results.
@@
Let $P_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. Let $B:\Sigma^{*}\rightarrow\{-1,+1\}$ be a behavior scoring function. The injections to all layers amounts to a change in the final hidden layer representation that is $q$ dependent, denoted by the vector $\delta r^{(L)}_{e}(q)$. Assume that the representations of aligned and misaligned answers w.r.t. $B$ are linearly separable, and that $\delta r^{(L)}_{e}(q)$ linearly classifies them with margin $\Delta$. Then, the behavior expectation of the model conditioned on the query $q$ satisfies:
@@
Where $B_{0}=B[P_{\theta}(\cdot|q)]$ is the behavior expectation without representation engineering and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
@@
As can be seen in the mathematical expression and in figure 2b for $B_{0}=-0.5$, this lower bound is a shifted hyperbolic tangent function w.r.t $r_{e}$. At $r_{e}=0$ the bound gives $B_{0}$, which is the unaltered model’s behavior. As $r_{e}$ is increased, the bound approaches $+1$, meaning the behavior asymptotically approaches $+1$. We see that for $B_{0}$ that is not too close to $-1$, the increase in behavior expectation is linear due to the hyperbolic tangent’s nature, while if it is very close to $-1$, $r_{e}$ is to be increased before seeing the linear effect. Thus for behaviors on which the model is negative but also has a small tendency for positive answers, the linear effect should be felt near $r_{e}=0$. In section 4, we present our numerical estimation $\Delta\lambda$ in the range $0.1-3$, both based on the linear classifier condition and direct alignment measurement. For proof see appendix section B.
@@
Figure 2: (a) The change to the last hidden layer due to vector injections from previous layers classifies positive and negative answer representations. (b) Plot of the upper bound on behavior expectation in theorem 1
Note that during decoding, the change in all layers of the model $R_{e}=\{r_{e}^{(1)},...,r_{e}^{(L-1)}\}$, amounts to a change to the final hidden layer’s representation, $r_{\theta,r_{e}}^{(L)}-r_{\theta}^{(L)}$, where $r_{\theta,r_{e}}^{(L)}$ is the final hidden layer representation that incorporates all the previous hidden layer changes and $r_{\theta}^{(L)}$ is the original representation. The complexity of the multi-layer editing is incorporated into the types of changes observed in the final hidden layer.
@@
In contrast to Wolf et al. (2023), that has a framework centralized on using prompts to misalign frozen models, i.e. models whose weights and representations are not changed after training, here the model is not frozen due to representation engineering, and accordingly a different result is obtained on guaranteeing an aligned response – for any adversarial attack, using large enough norms with representation engineering produces an aligned response if the learned injected representations accumulate to a good classifier of positive and negative answer representations in the final layer.
Corollary 1.
Let $\epsilon>0$, $P_{\theta}$ a language model and $q$ a prompt that induces negative behavior $B[P_{\theta}(\cdot|q)]<\gamma<0$ without representation engineering. Under the conditions of theorem 1, using an injected vector norm of $r_{e}>\frac{1}{\Delta\lambda}(arctanh(1-\epsilon)-arctanh(\gamma))$ leads to behavior expectation $B[P_{\theta,r_{e}}(\cdot|q)]>1-\epsilon$.
This can be extended to multi-token answers, by enforcing the above result on each decoding step of the generated answer, as explained in appendix I. The binary behavior score can also be extended beyond binary, as explained in appendix H.
@@
Now, we shall bound from above the helpfulness of the model as a function of representation engineering. We formally bound the probability of producing correct answers to queries where correctness is well-defined. Yet, even for queries where this is not the case, the bound can still be relevant, as it quantifies the model’s deviation from its original distribution due to representation engineering. Hence if the model was initially helpful on a task, a random deviation to its probability distribution is expected to decrease model performance proportionally to the size of the deviation.
@@
Intuitively, editing the model’s representation in a specific direction adds random noise to other latent concepts of the model, causing a degradation in its other capabilities. This is introduced in our framework through the resulting change to the final hidden layer $\delta r_{e}(q)=r_{\theta,r_{e}}^{(L)}-r_{\theta}^{(L)}$, we will assume its direction $\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|}$ contains random projections w.r.t. latent representations of correct and incorrect answers, which creates noise in the model’s distribution. The noise is expected to be random on the highest probability tokens, since they answer a query that is unrelated to the behavior being enhanced (intuitively depicted in figure 3a). We verify this empirically in appendix A.3. Thus, we assume random noise on the top $T$ tokens making up a large probability mass of the answer distribution, $1-\epsilon$, (e.g. $T\sim 10$ typically makes $\epsilon\sim 0.1$), and do not make assumptions on the rest of the vocabulary. The following theorem formally states this.
@@
Where $P_{0}=P_{\theta,r_{e}=0}(\cdot|q)$ is the probability of answering correctly without representation engineering, $T$ is the number of tokens making $1-\epsilon$ of the probability mass and $\alpha,\beta>0$ that depend on the query. $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.

### v5

# Tradeoffs Between Alignment and Helpfulness in Language Models with Steering Methods

@@ Abstract
Language model alignment has become an important component of AI safety, allowing safe interactions between humans and language models, by enhancing desired behaviors and inhibiting undesired ones. It is often done by tuning the model or inserting preset aligning prompts. Recently, steering methods, such as representation engineering and feature steering and activation steering, methods which alters the model’s behavior via changing its representations post-training, were shown to be effective in aligning LLMs. Steering methods yield gains in alignment oriented tasks such as resistance to adversarial attacks and reduction of social biases, but were also shown to cause a decrease in the ability of the model to perform basic tasks. In this paper we study the tradeoff between the increase in alignment and decrease in helpfulness of the model. We propose a theoretical framework which provides bounds for these two quantities, and demonstrate their relevance empirically. First, we find that under the conditions of our framework, alignment can be guaranteed with steering methods, and at the same time that helpfulness is harmed in the process. Second, we show that helpfulness is harmed quadratically with the norm of the injected steering vectors, while the alignment increases linearly with it, indicating a regime in which it is efficient to use representation engineering. We validate our findings empirically, and chart the boundaries to the usefulness of these methods for alignment. Our code is available at https://github.com/dorin133/REPE_alignment_helpfulness_tradeoff.
@@ 1 Introduction
Advancements in large language model (LLM) development over the last few years have given LLMs a variety of abilities that allow them to serve as general purpose assistants in a wide range of tasks, such as broad-scoped question answering, writing assistance, teaching, and more  (Radford et al., 2019; Devlin et al., 2019; Brown et al., 2020; Schulman et al., 2023; OpenAI, 2023; Bubeck et al., 2023; Nori et al., 2023; West, 2023; Park et al., 2023). The vast use of LLMs for such purposes has raised concerns due to the harm they can cause their users, such as serving fake information (Lin et al., 2022; Weidinger et al., 2022), behaving offensively, feeding social biases (Hutchinson et al., 2020; Venkit et al., 2022; Weidinger et al., 2022), or encouraging problematic behaviors by users Roose (2023); Atillah (2023). Alignment is often the term given for the process of removing these undesired behaviors (Yudkowsky, 2001; Taylor et al., 2016; Amodei et al., 2016; Shalev-Shwartz et al., 2020; Hendrycks et al., 2021; Pan et al., 2022; Ngo, 2022).
@@
There are several different approaches to performing alignment in LLMs, such as including aligning prompts (Askell et al., 2021; Rae et al., 2021) which was shown to improve alignment and decrease toxicity in LLMs, and the procedure of reinforcement learning from human feedback (RLHF) which trains language models to be helpful and harmless (Bai et al., 2022). Though effective to an extent, these approaches are still dangerously frail, as several works have shown that adversarial prompts can trigger negative behaviors in LLMs  Wallace et al. (2019); Yu & Sagae (2021); Xu et al. (2021); Subhash (2023); Zou et al. (2023b). The work of Wolf et al. (2023) provides a theoretical framework which shows that frozen LLMs can be misaligned with sufficiently long prompts.
Recently, new alignment methods were proposed, revolving around altering model weights at inference time, which control the model at the internal representations level by adding tailored vectors to the hidden layer’s representations. The appeal of such methods is that enhancing concepts through finetuning is expensive and not always efficient for small changes, while inference time steering requires only inference compute and allows to specialize the model to the user’s needs. Prominent methods include representation engineering (Zou et al., 2023a) and activation steering (Turner et al., 2023), in which directions in the model’s latent space controlling certain behaviors are extracted by contrasting hidden representations in which opposing behaviors are exhibited, as well as feature steering, by Anthropic (Templeton, 2024), in which steering vectors are obtained via the use of variational auto-encoders (VAEs), and demonstrate SOTA models such as Claude 3 Sonnet can be effectively steered by this method. While the methods differ in their approach for obtaining the steering vectors, the underlying principle of injecting the vectors into the model is simlar.
@@
Since then, there has been an increasing body of work using these methods. Zou et al. (2023a) demonstrated experimentally that the procedure can significantly improve alignment, e.g., in resistance to adversarial attacks, with reduction from 50% success of adversarial attacks to less than 15%, and truthfulness enhancement, with a relative increase of over 50%, though at the cost of somewhat reducing the helpfulness of the model. Wang et al. (2024b) use extracted safety vectors for inference time alignment for harmlessness, reducing jailbreaking success rate from over 30% with prompting and over 10% in supervised fine tuning to below one percent. Similar methods have also been used by Jorgensen et al. (2023); Leong et al. (2023); Liu et al. (2023); Turner et al. (2023) to improve alignment and reduce toxicity. Wang et al. (2024a) uses a method of editing model parameters that maximize the difference between toxic and untoxic responses to detoxify it. Wei et al. (2024) find sparse regions in parameter space that affect alignment brittleness, to be removed for better alignment. Marks et al. (2024) interpret causal graphs in language models and edit them to improve behaviors. van der Weij et al. (2024) extend activation steering to multiple behaviors. To improve low rank finetuning, Wu et al. (2024) utilize a procedure of tuning representations directly to substantially reduce the trainable parameters of finetuning compared to LoRA. Xu et al. (2024); Li et al. (2024) use concept activation vectors to jailbreak, they also observe that concepts that activate different behaviors are linearly separable. Zhang et al. (2024) remove hallucinations by editing truthfulness concepts. Additionally, the method scales to SOTA models, such as Claude 3 Sonnet (Templeton, 2024), using a similar method of sparse auto encoders, which extracts interpretable features from the model that can be used to manipulate the model through steering. There are also known limitations to editing representations - Yan et al. (2024) study limitations of model editing methods for social debiasing, and Elazar et al. (2021) empirically demonstrate how projecting out supervised linear probe directions can reduce performance on selected tasks.
@@
Understanding the tradeoff between model helpfulness and alignment is important for designing safe yet useful LLM systems. Previous empirical works have shown tradeoffs between quality and diversity and between helpfulness and safety in LLMs due to instruct finetuning (Florian et al., 2024; Bianchi et al., 2023; Röttger et al., 2023), and reduction in performance due to watermarking (Ajith et al., 2023). In this work we aim to shed light on the benefits and limitations of steering for LLM alignment, i.e., how much does alignment improve with this method and what is the cost in terms of the model’s abilities. We approach this question theoretically at first, and then provide empirical evidence for the validity of our theory.
@@
Figure 1: Effect of steering on helpfulness and alignment. Our main results show that alignment can improve at the cost of helpfulness. Moreover, we show that for small representation engineering norms the helpfulness decreases quadratically while the alignment increase is linear, so there is a regime in which representation engineering can be cost-effective.
@@
In sections 2 and 3, we set up our theoretical framework and present our theoretical results respectively. We find that steering increases alignment linearly with the steering vector norm (theorem 1), while the helpfulness of the model, defined as the probability of answering general queries correctly, decreases quadratically with the vector norm (theorem 2). Consequently, alignment can be guaranteed with large enough vector injections, though at the cost of reducing the model’s helpfulness. Conversely, when injecting vectors of small norms, the improvement of alignment is initially faster than the decrease in helpfulness, indicating a regime where steering is effective, allowing for inference time alignment while maintaining the model’s helpful capabilities. See figure 1 for an illustration of this intuition.
@@
In section 4 we explore the validity of our assumptions and results in an experimental setting with representation engineering: We calculate alignment, as defined by the theoretical framework, as a function of representation engineered vector norms corresponding to the desired behaviors and find that it increases as predicted by theorem 1. This is done by aligning with representation engineering an unaligned (pretrained) model with respect to desired behaviors (“harmless", “not-racist"), and misaligning an aligned (RLHF) model to undesired behaviors (“harmful", “racist"). Then, we calculate the helpfulness of the model, quantified by its question answering abilities over different knowledge domains and coding capabilities, with the same aligning vectors, and find that the decay with increased vector norm described in theorem 2 is manifested. Together, the results correspond to the intuitive illustration in fig. 1. Complementary experimental results by Anthropic showed similar empirical trends for alignment and helpfulness in the use of feature steering with vectors extracted from VAEs on Claude 3 Sonnet (Durmus et al., ).
@@ 2 Preliminaries
We denote $P_{\theta}(\cdot|s)$ as the next token probability distribution of a model with parameters $\theta$, when conditioned on the prompt $s$. The model is composed of $L$ layers, $r_{\theta}^{l}$ is the $l$’th hidden state representation of the model. The next token prediction of a model is parametrized as:
@@
Where $r^{(L)}_{\theta}(s)$ is the final hidden layer’s representation of the prompt $s$ and $U$ is an unembedding matrix from the hidden state to a vocabulary of tokens, a standard parametrization for SOTA LLMs. Denote a steered model by vectors, $R_{e}=(r_{e}^{(l=1)},...,r_{e}^{(l=L)})$, as $P_{\theta,r_{e}}$. Steering is performed at each layer by adding the corresponding vector to the hidden layer:
@@
Additionally, we follow existing methods for steering and provide a uniform norm for all the injected vectors $|r_{e}^{(l)}|=|r_{e}|$, which are initially prepared with norm $1$, and when injected to the model, are multiplied by the coefficient $r_{e}$ which can be positive or negative, to tune the steering strength and direction. For layers that are not injected, $|r_{e}^{(l)}|=0$.
To quantify alignment, we use the behavior expectation definition of alignment as in Wolf et al. (2023), based on the expected score of model responses to a behavior scoring function. The behavior scoring function can measure honesty, safety or any other concept for which responses can be scored as positively or negatively aligned with respect to. We will use a binary scoring function, with labels ±$\pm 1$ for aligned/misaligned answers. The results can be extended to more complex behavior scoring function over $[-1,+1]$, to yield qualitatively similar results, as discussed appendix J:
@@
While $B$ is a binary function, the behavior expectation is in the range $[-1,+1]$, reflecting cases where a model has probability for both aligned and misaligned responses. In theorem 1 we will prove that steering is an effective alignment method by lower bounding the behavior expectation. Notice that high probability of outputting a positive/negative response gives a positive/negative contribution to the behavior expectation, thus the sign and absolute value of behavior expectation measures the alignment of a model w.r.t. the given behavior.
The model’s helpfulness can be quantified as its ability to produce useful answers to user’s queries (knowledge questions, code generation, summarization, etc.). In order to theoretically analyze helpfulness, we focus on queries where correctness can be defined, such as knowledge based question answering (see figure 1 for an example) and code generation. This can be measured as the likelihood of outputting a correct answer to a query:
@@
Where $P_{\theta}(a_{correct}|q)$ is the model’s probability of outputting the correct answer $a$ to the query $q$. By this definition, the helpfulness is in the range $[0,1]$, in order to quantify the general capabilities of the model when steering vectors are injected into it. For queries where correctness is not defined, the bounds we derive are expected to still be meaningful as they also describe the rate of the model’s deviation from its original distribution due to steering.
The rational behind this quantification of alignment and helpfulness is to measure how aligning the model w.r.t. a concept through steering affects its ability to perform other tasks. Ideally, a model that interacts with a user should be both aligned and helpful, meaning its response is appropriate w.r.t. a desired behavior (quantified by a positive behavior expectation) and also useful (high probability of giving a correct answer to general purpose queries). In the next section, we will provide results on alignment and helpfulness under the use of steering, based on the model’s next token prediction, which provides simple analytical forms for alignment and helpfulness. In appendix K, we extend the results for multi-token answers, which yields qualitatively similar results, with somewhat more complex form.
@@ 3 Main Results
We will show that steering improves alignment and harms helpfulness, yet a "moderate" use of steering can yield a model that is good for both. Theorem 1 shows that behavior expectation is bounded from below by a hyperbolic tangent function, such that it approaches $+1$ for increasing size of injected vectors and increases linearly within a bounded range. This in principle allows to sample an aligned response for any adversarial attack (corollary 1), demonstrating the power of representation engineering as an alignment technique. Theorem 2 shows that the helpfulness is maximized in the vicinity of norm zero injected vectors (i.e., no representation engineering) and that as the norm is increased, helpfulness decays. The assumptions used to prove the theorems are presented formally in appendix A.
The following statement quantifies how alignment is improved by steering. It assumes the injected vectors in all layers accumulate to a change in the last hidden layer representation that classifies positive and negative behavior answers to the query, as depicted in figure 2a. This is assumed due to the popular choice in representation engineering to use steering vectors $\{r_{e}^{(l)}\}$, that are themselves classifiers for positive and negative representations on the intermediate layers, due to being learned from contrasting positive and negative behavior representations for different queries. For example, mean centering, $r_{e}^{(l)}=\mathbb{E}_{good,bad}[r_{good}^{(l)}-r_{bad}^{(l)}]$ (Jorgensen et al. (2023)), or PCA, $r_{e}^{(l)}=\text{argmax}_{v:||v||=1}[\mathbb{E}_{good,bad}|\langle v,r_{good}^{(l)}-r_{bad}^{(l)}\rangle|^{2}]$ (Zou et al. (2023a)), such that they form linear classifiers for the intermediate layers due to the positive/negative inner product with positive/negative answer representations. Notably, in Xu et al. (2024) it is shown empirically that such concept classes in latent space are linearly separable. We discuss this assumption further in A and provide empirical evidence. Furthermore, the classification condition can be softened to an imperfect classifier, as discussed in appendix A and shown in appendix in H, to yield similar results.
@@
Let $P_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. Let $B:\Sigma^{*}\rightarrow\{-1,+1\}$ be a behavior scoring function. The injections to all layers amounts to a change in the final hidden layer representation that is $q$ dependent, denoted by the vector $\delta r^{(L)}_{e}(q)$. Assume the representations of aligned and misaligned answers w.r.t. $B$ are linearly separable, and $\delta r^{(L)}_{e}(q)$ linearly classifies them with margin $\Delta$. Then, the behavior expectation of the model conditioned on the query $q$ satisfies:
@@
Where $B_{0}=B[P_{\theta}(\cdot|q)]$ is the behavior expectation without steering and $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
@@
As can be seen in the mathematical expression and in figure 2b for $B_{0}=-0.5$, this lower bound is a shifted hyperbolic tangent function w.r.t $r_{e}$. At $r_{e}=0$ the bound gives $B_{0}$, which is the unaltered model’s behavior. As $r_{e}$ is increased, the bound approaches $+1$, meaning the behavior asymptotically approaches $+1$. We see that for $B_{0}$ that is not too close to $-1$, the increase in behavior expectation is linear due to the hyperbolic tangent’s nature, while if it is very close to $-1$, $r_{e}$ is to be increased before seeing the linear effect. Thus for behaviors on which the model is negative but has a small tendency for positive answers, the linear effect should be felt near $r_{e}=0$. In section 4, we present our numerical estimation $\Delta\lambda$ in the range $0.1-3$, both based on the linear classifier condition and direct alignment measurement. For proof see appendix section B.
@@
Figure 2: (a) The change to the last hidden layer due to vector injections from previous layers classifies positive and negative answer representations. (b) Plot of the upper bound on behavior expectation in theorem 1.
@@
This can be extended to multi-token answers, by enforcing the above result on each decoding step of the generated answer, as explained in appendix K. The binary behavior score can also be extended beyond binary, as explained in appendix J. In contrast to Wolf et al. (2023), whose framework is centralized on using prompts to misalign frozen models, i.e.  whose weights and representations are not changed after training, here the model is not frozen due to steering, and accordingly a different result is obtained on guaranteeing an aligned response – for any adversarial attack, using steering with large enough norms produces an aligned response if the learned steering vectors accumulate to a good classifier of positive and negative answer representations in the final layer. We formalize this in appendix D.
@@
Now, we shall bound from above the helpfulness of the model as a function of steering. We formally bound the probability of producing correct answers to queries where correctness is well defined. Yet, even when this is not the case, the bound can still be relevant, as it quantifies the model’s deviation from its original distribution due to steering. Hence if the model was initially helpful on a task, a random deviation to its probability distribution is expected to decrease model performance proportionally to the size of the deviation.
@@
Intuitively, editing the model’s representation in a specific direction adds random noise to other latent concepts of the model, causing a degradation in its other capabilities. This is introduced in our framework through the resulting change to the final hidden layer $\delta r_{e}(q)=r_{\theta,r_{e}}^{(L)}-r_{\theta}^{(L)}$, we will assume its direction $\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|}$ contains random projections w.r.t. latent representations of correct and incorrect answers, which creates noise in the model’s distribution. The noise is expected to be random on the highest probability tokens, when answering a query that is unrelated to the behavior being enhanced (intuitively depicted in figure 3a). We verify this empirically in appendix A.3. Thus, we assume random noise on the top $T$ tokens making up a large probability mass of the answer distribution, $1-\epsilon$, (e.g. $T\sim 10$ typically makes $\epsilon\sim 0.1$), and do not make assumptions on the rest of the vocabulary. The following theorem formally states this.
@@
Where $P_{0}=P_{\theta,r_{e}=0}(\cdot|q)$ is the probability of answering correctly without steering, $T$ is the number of tokens making $1-\epsilon$ of the probability mass and $\alpha,\beta>0$ that depend on the query. $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.